OpenStudy (anonymous):

How to find the arclength of (1/16)x^4 + (1/2)x^-2 on the interval of [1,2]?

OpenStudy (anonymous):

You have the function, now apply the general form of your arc length equation: \[L = \int\limits \ { \sqrt{1 + (\frac{dy}{dx})^2} } dx\]

OpenStudy (anonymous):

The derivative of the function shouldn't be too difficult to find, so once you have that, square it, add one, take the square root of the entire expression and integrate w.r.t. x between 1 and 2 to get the value for the length.